Given that 𝑓(π‘₯) = 3π‘₯ 2 βˆ’ 4π‘₯ + 7, use the definition of the derivative to find 𝑓 β€² (π‘₯).


1
Expert's answer
2021-06-08T17:30:31-0400

Given that "\ud835\udc53(\ud835\udc65) = 3\ud835\udc65^2 \u2212 4\ud835\udc65 + 7" , let us find "\ud835\udc53'(\ud835\udc65)":


"f'(x)=\lim\limits_{\Delta x\to 0}\frac{f(x+\Delta x)-f(x)}{\Delta x}=
\lim\limits_{\Delta x\to 0}\frac{3(x+\Delta x)^2-4(x+\Delta x)+7-(3x^3-4x+7)}{\Delta x}=
\lim\limits_{\Delta x\to 0}\frac{3x^3+6x\Delta x+3(\Delta x)^2-4x-4\Delta x+7-3x^2+4x-7}{\Delta x}=
\lim\limits_{\Delta x\to 0}\frac{6x\Delta x+3(\Delta x)^2-4\Delta x}{\Delta x}=
\lim\limits_{\Delta x\to 0}(6x+3\Delta x-4)=6x-4."